### v1

@@ 3.4 Synthetic data tests
Figure 3: Results for the synthetic data tests described in Section D.2. The ground truth “scenarios” are labeled on the $x$-axis. Here, we show the mean and central 90% of continuous rank probability scores [CRPS; 49] and root mean square error (RMSE) across 100 synthetic data sets for three competing methods relative to fitting a Gaussian process with a stationary Matérn kernel (M1). Scores are tallied for each of the data generating mechanisms describe in Table D.1.
@@
In order to ease comparison, we show CRPS and RMSE for M2-M4 relative to corresponding quantities from M1: relative scores less than one indicate that RMSE and CRPS are smaller (and therefore better) compared with a stationary GP. Results are shown in Figure 3, where we show the average (relative) scores and the 5%-95% range over the $100$ draws. When the underlying data are non-sparse and stationary (Scenario 1), i.e., M1 is the “correct” model, methods M2-M4 all perform as well as M1 in terms of both CRPS and RMSE. This makes sense for the Vecchia approximation, since much of the underlying theory is derived for traditional (i.e., non-sparse) stationary kernels; it is reassuring that our kernel (which is nonstationary by definition for both M3 and M4) performs well for stationary data. When the ground truth is sparse and stationary (Scenario 2), M3 and M4 perform similarly–albeit slightly better–than M1, while the Vecchia approximation (M2) performs slightly worse than M1. However, when the underlying true state involves nonstationarity (Scenarios 3 and 4), the nonstationary kernels (M3 and M4) yield a noticeable improvement over a stationary GP, particularly in terms of CRPS. In contrast, M2 is noticeably worse for these scenarios, which suggests that the standard configuration of the sparse general Vecchia approximation is sub-optimal for non-stationary data. The advantage of our kernel is particularly noticeable when the underlying ground truth is nonstationary and sparse.
@@
In summary, these comparisons verify that our kernel performs as well or better than more traditional kernels within a GP across a variety of ground truth data-generating mechanisms. Since each of the data-generating mechanisms considered (different combinations of stationary/nonstationary and sparse/nonsparse) are special cases of our kernel (as described in Section 2.3), this furthermore verifies that our kernel can flexibly learn the true underlying state of a given data set.
4 Application: space-time prediction of daily temperature
@@
Probabilistic machine learning methods like GPs are well-suited for conducting spatio-temporal prediction of in situ measurements of daily maximum temperature. While these data are the gold standard in the Earth sciences, their irregular sampling presents a challenge for climate and atmospheric scientists, who often require data on regular space-time units. To address this need, there are a large number of so-called gridded data products that use (in many cases) naïve statistical methods to conduct spatio-temporal interpolation. For example, the Livneh et al. (2015) [50, 51] data product (henceforth L15) provides daily estimates of surface temperature on a regular grid over the United States from 1950 to present. The L15 data are generated using inverse-distance weighted averaging [52] separately for each day’s measurements before applying a post-hoc correction for orography; no measures of uncertainty quantification are provided. GPs provide many value-added benefits to conducting space-time prediction for temperature measurements relative to the L15 methodology: data-driven (not data-agnostic) weights for spatial averaging; nonstationary weights for interpolation; the ability to account for multiple-day autocorrelation in temperature records; a flexible prior mean function to capture known spatio-temporal structure and systematically estimate relationships between daily temperature and orography; and most importantly uncertainty quantification.
4.1 Spatial modeling of data time slices
@@
We first apply the kernel $C_{z}$ and the Bayesian model to time slices of daily maximum temperature (C) from the GHCN-D database [53], considering stations that lie within the contiguous United States (CONUS). These measurements are the same input data used to construct the L15 data product. We identify a set of $N_{\text{space}}=1174$ GHCN-D stations with high-quality records (see Figure F.1a.) and randomly select 60 days for training: one from each month over 2001-2005. In this case, the input space collapses down to two-dimensional coordinates, longitude and latitude, such that $\mathcal{X}\subset\mathbb{R}^{2}$. For test data, we use measurements from $N_{\text{space, test}}=1444$ records not included in the training records.
@@
We refer the interested reader to Appendix F.1 for full details on the prior mean function and specific kernel used for the time slices. In summary, the prior mean is a linear function of orographic features and distance to the nearest coastlines to account for known relationships between these features and daily temperature. The core kernel is the locally anisotropic nonstationary kernel defined in [45] wherein the signal variance and anisotropy length scales are suitably transformed linear functions of elevation and distance to the coast. For the sparsity-inducing kernel, following the guidance of Section D.1 we err on the side of choosing $n_{1}$ and $n_{2}$ to be larger than we might expect. We consider three configurations, each with 100 total bump functions: (1) $n_{1}=2$ and $n_{2}=50$; (2) $n_{1}=4$ and $n_{2}=25$; and (3) $n_{1}=10$ and $n_{2}=10$. Furthermore, we consider three upper limits on the Wendland truncation radii (5, 10, and 15 units of degrees longitude/latitude) to ensure that the resulting kernel has the chance to impose sparsity on the data. Ultimately, our specification results in five hyperparameters for the prior mean function, twelve hyperparameters for the core kernel, one hyperparameter for the noise variance, and $5033+n_{1}n_{2}(d+3)=3+100\times 5=$ 503 hyperparameters for the sparse kernel: in total there are $5+12+1+503=521$ hyperparameters for the GP applied to each time slice.
@@
Results. Before analyzing all 60 time slices, we first looked at the January 14, 2001 time slice (training and test data shown in Figure 5) to determine an optimal configuration of $n_{1}$ and $n_{2}$ as well as the Wendland length scale upper bound. Three $n_{1}\times n_{2}$ configurations ($10\times 10$, $4\times 25$, $2\times 50$) and three upper bound thresholds result in nine different sparse-inducing GP models. For each model, the MCMC was run for 5,000 iterations, with the first 4,001 discarded as burn-in; convergence of the MCMC was assessed graphically with no issues detected. All subsequent calculations were based on the posterior median of the hyperparameters. The log of the marginal likelihood of the training data, conditioned on the posterior median, the resulting sparsity of the training data covariance, as well as out-of-sample root mean square error (RMSE) and the continuous rank probability score [CRPS; 49] for the test data are shown in Figure 4 for each of the nine models. For comparison, Figure 4 also shows corresponding quantities from a similar Gaussian process model with the same prior mean and core kernel but without the sparse kernel (dashed lines). A few interesting results emerge from the different models considered: first, as expected, the sparsity increases monotonically with the Wendland upper bound; however, unexpectedly, there is a clear trend in the sparsity across the $n_{1}\times n_{2}$ configurations, wherein larger values of $n_{1}$ lead to a smaller fraction of nonzero elements in the covariance. The in-sample log-likelihood and out-of-sample RMSE/CRPS all greatly prefer longer Wendland length-scales (upper bound of 10 or 15), and from a testing standpoint the sparse nonstationary GPs perform as well as the nonsparse nonstationary GP (with about $1/3$ the sparsity). As a compromise between good out-of-sample performance and sparsity, we conclude that the $n_{1}=4$ and $n_{2}=25$ configuration with upper bound 10 is the “optimal” configuration to use for in situ measurements of daily maximum temperature.
[…]
@@
Figure 5: Training and test data for the January 14, 2001 time slice of daily maximum temperature (panels a. and b., respectively). Panels (c) and (d) show the spatial distribution of interpolated daily maximum temperature for both the inverse-distance weighting and our method. The inset text in panels (c) and (d) shows the RMSE for the test data using each method.
@@
Using the optimal configuration, we show visualizations of the prior mean and covariance hyperparameters for the January 14, 2001 time slice in Figure G.2. Since the prior mean function and core kernel parameters are specified as functions of physical, topographical covariates, we can visualize the implied surfaces for the mean function, core kernel signal variance, core kernel anisotropy length-scales, and implied correlation between a set of reference locations and all other points in a $0.25^{\circ}\times 0.25^{\circ}$ longitude-latitude grid over the United States. These quantities (see Figure G.2) provide useful scientific information regarding the first- and second-order properties of daily temperatures, and how these vary according to topography. We can furthermore see how the bump functions impose regions of distance-unrelated correlations across the United States. Finally, we compare predictions for a $0.25^{\circ}\times 0.25^{\circ}$ longitude-latitude grid in Figure 5 for the L15 inverse-distance methodology with corresponding GP-based estimates; see Figure 5(c) and (d). Our methodology yields a clear improvement over L15: both with respect to reduced root mean square error for the test data and also for interpolation of daily maximum temperature in mountainous terrain. For example, note that the inverse-distance weighting completely fails to characterize the topographic variability in, e.g., Nevada and California. Furthermore, it is clear that the L15 method “under-smooths” the data in the eastern United States, particularly along the Gulf Coast.
@@
Now stepping back to consider all 60 time slices, we show corresponding in-sample log-likelihood, sparsity, and out-of-sample metrics in Table 1. For reference, we also summarize RMSE for the L15 method. It is once again clear that a sparse nonstationary GP vastly outperforms L15: on average, the errors are roughly $100(3.00-2.54)/3.00=15.3\%$ smaller (furthermore, the nonstationary GP errors are uniformly smaller than the L15 errors across all time slices). As before, the $n_{1}=4$ and $n_{2}=25$ configuration results in about 37% sparsity, on average.
@@
Table 1: Out-of-sample root mean square error (RMSE) and continuous rank probability score (CRPS) as well as in-sample log-likelihood and sparsity (defined as the fraction of the covariance that is nonzero) aggregated over all 60 time-slices. For comparison, we show RMSE using the inverse-distance weighting methodology of L15. The columns show the average, minimim, and maximum taken over the 60 time-slices.
Method Metric Average Minimum Maximum
L15 RMSE (test data; C) 3.00 2.38 4.22
Nonstationary, sparse GP RMSE (test data; C) 2.54 1.89 4.16
CRPS (test data; no units) 1.50 1.11 2.92
Log-likelihood (training data; no units) -2673.77 -3280.88 -2303.74
Sparsity (fraction of nonzero covariance) 0.37 0.31 0.46
@@ 4.2 Spatio-temporal modeling
We next apply the Bayesian model proposed in Eq. 3.1 to spatio-temporal measurements from the $N_{\text{train}}=1174$ stations over 2001-2005 ($N_{\text{time}}=1826$ days), such that the input space is $\mathcal{X}=\mathcal{S}\times\mathcal{T}$ (two-dimensional space and one dimension for time), i.e., $d=3$. To make the calculation more feasible within the allotted time frame, we pick every second data point which yields a total of $N=1,038,723$ non-missing daily measurements for training. We again refer the interested reader to Appendix F.2 for full details on the prior mean function and specific kernel used for the spatio-temporal measurements. In summary, in light of the fact that we are modeling high-dimensional space-time data and our kernel can discover and capitalize upon sparsity, it is to our benefit to propose a prior mean function that explains as much of the known spatial and temporal structure in the data as possible. The most prominent structure in a data set involving temperature is the annual cycle, wherein daily maximum temperatures in the Northern Hemisphere are generally lowest in the winter and highest in the summer. As in Section 4.1, there is an underlying geospatial climatology to measurements of temperature including an implicit zonal structure wherein spatial locations farther from the equator are generally cooler, a strong influence from orographic variability, and the influence of coastlines.
@@
To account for each of these features, we propose a flexible prior mean function using natural cubic splines with spatially-varying coefficients defined by thin-plate spline basis functions and orography. After some trial and error and in order to maintain a compromise between sufficient flexibility and overfitting, we end up with $1899$ hyperparameters for the prior mean: one for every $\approx 550$ data points. Using ordinary least squares, the prior mean function explains 94.1% of the variability in the daily temperature data and removes meaningful autocorrelations beyond $\approx 5$ days (see Figure F.1). In other words, the variability explained by a flexible prior mean function “increases” the sparsity that can be leveraged by the kernel. The space-time core kernel is similar to the one in Equation F.1, albeit generalized to a three-dimensional input space, i.e., $C_{\text{core}}({\bf x},{\bf x}^{\prime})$, with stationarity in the time direction. It is important to note that this specification yields a non-separable space-time kernel, i.e., $C_{\text{core}}({\bf x},{\bf x}^{\prime})\neq C({\bf s},{\bf s}^{\prime})\times C(t,t^{\prime})$. For the sparse kernel, the Wendland component is again anisotropic with coordinate-specific length scales (and coordinate-specific truncation limits); following Section 4.1 we use 10 units for the spatial truncation limits and 4 days for the temporal truncation limits based on the fact that the prior mean function explains autocorrelation over longer time lags. We again use 100 bump functions with $n_{1}=4$ and $n_{2}=25$; however, based on the fact that the prior mean explains so much of the temporal variability, we restrict the bump functions to live in the spatial domain only. This yields $521+1+1=523$ hyperparameters for the kernel; combined with the prior mean hyperparameters, we have $2,422$ hyperparameters to be learned in training with MCMC.
@@
To summarize the results of this analysis, we selected seven test sites from across the United States and assessed the time-evolving behavior of daily maximum temperatures over 2001-2005; see Figure 6 (the first two years are omitted for visibility of results). The training sites with test locations are shown on the top left, with the test data and posterior predictions shown in each thumbnail. The most obvious feature of the predictions is the annual cycle, which is well-characterized by our prior mean function. However, note that the specific shape of the annual cycle is flexibly modeled across sites: temperatures in Houston, TX (top right) increase sharply in late winter and decrease sharply in late fall, while temperatures in Portland, OR (middle left) increase and decrease much more slowly between the cold and warm season. The GP predictions maintain the large day-to-day variations present in daily maximum temperatures, which is reflected in the data. Also note that the borrowing of strength over space and time allows the GP to learn the time-varying structure even when there are large chunks of missing data in the traning data, e.g., in the first half of 2003 at the Houston, TX gauged location. Finally, the three-sigma GP uncertainties, represented by the red shaded band in Figure 6, are clearly nonstationary as well: note the larger magnitude of uncertainties at, e.g., Houston, TX compared to the smaller uncertainties for the Berkeley, CA site.

### v3

@@ 4.2 Comparison with competing methods
Figure 5: Results for the synthetic data tests described in Section 4.2. The ground truth “scenarios” are denoted with different colors, while the different signal-to-noise (SNR) considered are shown on the $x$-axis. We show the mean and central 90% of continuous rank probability scores [CRPS; 52] across 50 synthetic data sets for five competing methods relative to fitting a Gaussian process with a stationary Matérn kernel (M1); relative scores less than 100% indicate better performance relative to M1. Scores are tallied for each of the data generating mechanisms describe in Table 1.
Results are shown in Figure 5, where we show the average (relative) scores and the 5%-95% range over the $N_{\text{rep}}=50$ repeated draws. When the underlying data are non-sparse and stationary (Scenario 1), i.e., M1 is the “correct” model, all methods except covariance tapering perform nearly as well as M1. This makes sense for the Vecchia approximation, since much of the underlying theory is derived for traditional (i.e., non-sparse) stationary kernels; similarly, the predictive process has been shown to perform well in capturing large-scale features (as is present with a non-sparse kernel). For the same reason, it is unsurprising that covariance tapering performs worse than the other methods for non-sparse data, as in Scenario 1. It is reassuring that our kernel (which is nonstationary by definition for both M5 and M6) performs well for stationary data. A slight exception to this is when the SNR is low: in this case, M5 and M6 perform worse, on average, than M1. This too is unsurprising, because when the noise is large and the ground truth is smooth, a sparse kernel could overfit the noise. This has important implications for the sorts of data sets that may most benefit from our proposed kernel: since sparsity is a key feature of the kernel, it may not be best suited to very smooth functions.
@@
For Scenario 2, where the ground truth is sparse and stationary, M2 and M4 perform nearly as well as M1, while M3 performs significantly worse (indicating that the predictive process struggles to capture sparse structure, as has been observed elsewhere). M5 and M6 perform as well as M1 when the SNR is large, but start to show large increases in performance (on average, 10-15%) when the SNR is small. However, when the underlying true state involves nonstationarity (Scenarios 3 and 4), the nonstationary kernels (M5 and M6) yield a noticeable improvement over a stationary GP, particularly for low SNR. In contrast, M2 and M3 are noticeably worse for these scenarios, which suggests that the standard configuration of the sparse general Vecchia approximation and the predictive process is sub-optimal for non-stationary data. The advantage of our kernel is particularly noticeable when the underlying ground truth is nonstationary and sparse.
@@
In summary, these comparisons verify that our kernel performs as well or better than more traditional kernels within a GP across a variety of ground truth data-generating mechanisms and a range of SNR values. The improvements in performance are admittedly modest for SNR = 10 or 20, but for SNR = 5 the relative improvement for M5 and M6 vs. M1 is on 12-15% on average and as large as 50% for some synthetic data sets. In other words, for large noise, when the underlying data are sparse and/or nonstationary, the stationary Matérn fails to characterize the true function; however, even in these cases, M5 and M6 can capture the ground truth. This furthermore verifies that each of the data-generating mechanisms (different combinations of stationary/nonstationary and sparse/nonsparse) are special cases of our kernel (as described in Section 2.3); in other words, our kernel can learn the true underlying state of a given data set.
5 Application: space-time prediction of daily maximum temperature
@@
Probabilistic machine learning methods like GPs are well-suited for conducting spatio-temporal prediction of in situ measurements of daily maximum temperature. While these data are the gold standard in the Earth sciences, their irregular sampling presents a challenge for climate and atmospheric scientists, who often require data on regular space-time units. To address this need, there are a large number of so-called gridded data products that use (in many cases) naïve statistical methods to conduct spatio-temporal interpolation. For example, the Livneh et al. (2015) [55, 56] data product (henceforth L15) provides estimates of daily maximum surface temperature on a regular grid over the United States from 1950 to present. The L15 data are generated using inverse-distance weighted averaging [57] separately for each day’s measurements before applying a post-hoc correction for orography; no measures of uncertainty quantification are provided. GPs provide many value-added benefits to conducting space-time prediction for daily maximum temperature measurements relative to the L15 methodology: data-driven (not data-agnostic) weights for spatial averaging; nonstationary weights for interpolation; the ability to account for multiple-day autocorrelation in daily maximum temperature records; a flexible prior mean function to capture known spatio-temporal structure and systematically estimate relationships between daily maximum temperature and orography; and most importantly uncertainty quantification.
5.1 Spatial modeling of data from individual days
@@
We first apply the kernel $C_{z}$ and the Bayesian model to all measurements from a given day of daily maximum temperature (C) from the GHCN-D database [58], considering stations that lie within the contiguous United States (CONUS). These measurements are the same input data used to construct the L15 data product. We identify a set of $N_{\text{space}}=1174$ GHCN-D stations with high-quality records (see Figure E.1a.) and randomly select 60 days for training: one from each month over 2001-2005. In this case, the input space collapses down to two-dimensional coordinates, longitude and latitude, such that $\mathcal{X}\subset\mathbb{R}^{2}$. For test data, we use measurements from $N_{\text{space, test}}=1444$ records not included in the training records.
@@
We refer the interested reader to Appendix E.1 for full details on the prior mean function and specific kernel used for spatial modeling . In summary, in order to maintain a fair comparison with the L15 methodology, the prior mean is set to be a constant. The core kernel is the locally anisotropic nonstationary kernel defined in [46] wherein the signal variance and anisotropy length scales are suitably transformed linear functions of elevation . For the sparsity-inducing kernel, following the guidance of Section 4.1 we err on the side of choosing $n_{1}$ and $n_{2}$ to be larger than we might expect. We consider three configurations, each with 100 total bump functions: (1) $n_{1}=2$ and $n_{2}=50$; (2) $n_{1}=4$ and $n_{2}=25$; and (3) $n_{1}=10$ and $n_{2}=10$. Furthermore, we consider three upper limits on the Wendland truncation radii (5, 10, and 15 units of degrees longitude/latitude) to ensure that the resulting kernel has the chance to impose sparsity on the data. Ultimately, our specification results in one hyperparameter for the prior mean function, eight hyperparameters for the core kernel, one hyperparameter for the noise variance, and $3+n_{1}n_{2}(d+3)=3+100\times 5=503$ hyperparameters for the sparse kernel: in total there are $1+8+1+503=513$ hyperparameters for the GP applied to each day’s measurements .
@@
Results. Before analyzing all 60 days , we first looked at data from January 14, 2001 (training and test data shown in Figure 6) to determine an optimal configuration of $n_{1}$ and $n_{2}$ as well as the Wendland length scale upper bound. Three $n_{1}\times n_{2}$ configurations ($10\times 10$, $4\times 25$, $2\times 50$) and three upper bound thresholds result in nine different sparse-inducing GP models. For each model, the MCMC was run for 20,000 iterations, with the first 15,001 discarded as burn-in; convergence of the MCMC was assessed graphically with no issues detected (trace plots from the $4\times 25$ configuration are shown in Figures F.3 and F.4). All subsequent calculations were based on the posterior median of the hyperparameters. The log of the marginal likelihood of the training data, conditioned on the posterior median, the resulting sparsity of the training data covariance, as well as out-of-sample root mean square error (RMSE) and the continuous rank probability score [CRPS; 52] for the test data are shown in Supplemental Figure F.2 for each of the nine models. For comparison, Supplemental Figure F.2 also shows corresponding quantities from a similar Gaussian process model with the same prior mean and core kernel but without the sparse kernel (dashed lines). A few interesting results emerge from the different models considered: first, as expected, the sparsity increases monotonically with the Wendland upper bound; however, unexpectedly, there is a clear trend in the sparsity across the $n_{1}\times n_{2}$ configurations, wherein larger values of $n_{1}$ lead to a smaller fraction of nonzero elements in the covariance. The in-sample log-likelihood and out-of-sample RMSE/CRPS all greatly prefer longer Wendland length-scales (upper bound of 10 or 15), and from a testing standpoint the sparse nonstationary GPs perform as well as the nonsparse nonstationary GP (with about $1/3$ the sparsity). As a compromise between good out-of-sample performance and sparsity, we conclude that the $n_{1}=4$ and $n_{2}=25$ configuration with upper bound 10 is the “optimal” configuration to use for in situ measurements of daily maximum temperature.
@@
Figure 6: Training and test data for January 14, 2001 daily maximum temperature (panel a.). Panels (b) and (c) show the spatial distribution of interpolated daily maximum temperature for inverse-distance weighting and the various Gaussian process methods, while panel (d) shows the Gaussian process uncertainties (posterior standard deviation). The inset text in panels (b) and (c) shows the RMSE and CRPS for the test data using each method.
@@
Using the optimal configuration, we show visualizations of the covariance hyperparameters for the data from January 14, 2001 in Figure F.5. Since the core kernel parameters are specified as functions of physical, topographical covariates, we can visualize the implied surfaces for the core kernel signal variance, core kernel anisotropy length-scales, and implied correlation between a set of reference locations and all other points in a $0.25^{\circ}\times 0.25^{\circ}$ longitude-latitude grid over the United States. These quantities (see Figure F.5) provide useful scientific information regarding the second-order properties of daily maximum temperatures, and how these vary according to topography. We can furthermore see how the bump functions impose regions of distance-unrelated correlations across the United States. We note that small changes to the reference location can result in very different correlation structures across the rest of the domain. Since our proposed kernel is highly flexible and fully data-driven (with only a relatively small penalty for regularization), it is difficult to say whether these correlations are a property of the underlying data generation mechanism or an artifact related to this specific data sample.
Finally, we compare predictions for a $0.25^{\circ}\times 0.25^{\circ}$ longitude-latitude grid in Figure 6 for the L15 inverse-distance methodology with corresponding GP-based estimates; see Figure 6(c) and (d). In addition to our new methodology, we also show GP-based estimates from models M2 (Vecchia), M3 (predictive process), and M4 (covariance tapering), since the L15 methodology is a deterministic, non-model-based technique that does not account for covariates or uncertainty. Our methodology yields a clear improvement over L15 and the other GP predictions: both with respect to reduced root mean square error and CRPS for the test data, and also for interpolation of daily maximum temperature in mountainous terrain. For example, note that the inverse-distance weighting and competing GP methods completely fail to characterize the topographic variability in, e.g., Nevada and California. Furthermore, it is clear that the L15 method “under-smooths” the data in the eastern United States, particularly along the Gulf Coast. Our nonstationary, sparse GP also yields heterogeneous uncertainty estimates, reflecting the underlying differences in the degree of uncertainty in temperature measurements in flat versus topographically diverse regions.
@@
Now stepping back to consider all 60 days , we show out-of-sample prediction metrics for our proposed method in Table 2. For reference, we also summarize RMSE for inverse distance weighting used by the L15 method and the other GP-based methods. It is once again clear that a sparse nonstationary GP vastly outperforms L15: on average, the errors are roughly $100\%(3.00-2.62)/3.00=12.6\%$ smaller. Furthermore, our nonstationary GP outperforms the other GP-based methods considered here, providing the best RMSE and CRPS on average.
RMSE (test data; C) CRPS (test data; no units)
Method Average Minimum Maximum Average Minimum Maximum
Inverse distance weighting 3.00 2.38 4.22
Vecchia approximation 2.79 2.25 3.80 1.53 1.22 2.05
Predictive process 2.84 2.30 3.80 1.54 1.23 2.07
Covariance tapering 3.04 2.33 4.20 1.65 1.27 2.22
Nonstationary, sparse GP 2.62 2.02 3.74 1.50 1.09 2.15
@@
Table 2: Out-of-sample root mean square error (RMSE) and continuous rank probability score (CRPS) for inverse-distance weighting, three competing methods, and our new methodology. Metrics are aggregated over all 60 days, showing the average, minimum, and maximum taken over the 60 days. Note: CRPS cannot be calculated for inverse distance weighting because the methodology does not yield uncertainty measures.
@@ 5.2 Spatio-temporal modeling
We next apply the Bayesian model proposed in Eq. 3.1 to spatio-temporal measurements from the $N_{\text{train}}=1174$ stations over 2001-2005 ($N_{\text{time}}=1826$ days), such that the input space is $\mathcal{X}=\mathcal{S}\times\mathcal{T}$ (two-dimensional space and one dimension for time), i.e., $d=3$. To make the calculation more feasible within the allotted time frame, we pick every second data point which yields a total of $N=1,038,723$ non-missing daily measurements for training. We again refer the interested reader to Appendix E.2 for full details on the prior mean function and specific kernel used for the spatio-temporal measurements. In summary, in light of the fact that we are modeling high-dimensional space-time data and our kernel can discover and capitalize upon sparsity, it is to our benefit to propose a prior mean function that explains as much of the known spatial and temporal structure in the data as possible. The most prominent structure in a data set involving temperature is the annual cycle, wherein daily maximum temperatures in the Northern Hemisphere are generally lowest in the winter and highest in the summer. As in Section 5.1, there is an underlying geospatial climatology to measurements of daily maximum temperature including an implicit zonal structure wherein spatial locations farther from the equator are generally cooler, a strong influence from orographic variability, and the influence of coastlines.
@@
To account for each of these features, we propose a flexible prior mean function using natural cubic splines with spatially-varying coefficients defined by thin-plate spline basis functions and orography. In order to maintain a compromise between sufficient flexibility and overfitting, we end up with $1899$ hyperparameters for the prior mean: one for every $\approx 550$ data points. Using ordinary least squares, the prior mean function explains 94.1% of the variability in the daily maximum temperature data and removes meaningful autocorrelations beyond $\approx 5$ days (see Figure E.1). In other words, the variability explained by a flexible prior mean function “increases” the sparsity that can be leveraged by the kernel. The space-time core kernel is similar to the one in Equation E.1, albeit generalized to a three-dimensional input space, i.e., $C_{\text{core}}({\bf x},{\bf x}^{\prime})$, with stationarity in the time direction. It is important to note that this specification yields a non-separable space-time kernel, i.e., $C_{\text{core}}({\bf x},{\bf x}^{\prime})\neq C({\bf s},{\bf s}^{\prime})\times C(t,t^{\prime})$. For the sparse kernel, the Wendland component is again anisotropic with coordinate-specific length scales (and coordinate-specific truncation limits); following Section 5.1 we use 10 units for the spatial truncation limits and 4 days for the temporal truncation limits based on the fact that the prior mean function explains autocorrelation over longer time lags. We again use 100 bump functions with $n_{1}=4$ and $n_{2}=25$; however, based on the fact that the prior mean explains so much of the temporal variability, we restrict the bump functions to live in the spatial domain only. This yields $521+1+1=523$ hyperparameters for the kernel; combined with the prior mean hyperparameters, we have $2,422$ hyperparameters to be learned in training with MCMC.
@@
To summarize the results of this analysis, we selected seven test sites from across the United States (see the top left panel of Figure 8). We first explore the spatio-temporal correlations implied by our sparse, nonstationary kernel; for brevity, we focus on three of these sites and plot directional autocorrelation functions in Figure 7 (corresponding plots for the four other sites are shown in Figure F.6). Here, we define directional autocorrelation functions to be the time-evolving correlation for a pair of sites, i.e., $C_{y}\left([{\bf s},t],[{\bf s}^{\prime},t^{\prime}]\right)$ (normalized to show correlation instead of covariance), where $\bf s$} is one of the reference cities and $\bf s$$^{\prime}$ is a point to the north, south, east, or west of the reference; since our kernel is stationary in the time domain we set $t=0$ and $t^{\prime}\in[0,5]$. Figure 7(a) also shows the activated bump functions (i.e., those with $a_{ij}=1$) and their size; recall that the bump functions are only defined in the spatial domain. The three selected sites each have unique correlation properties. The site of interest at Denver, CO is located within an activated bump function, and hence its autocorrelation function decays to zero but is never exactly zero. The same is true for the autocorrelation with points to the south and east, since these points share at least one bump function with Denver, CO. However, points to the north and west do not share a bump function, which means that their autocorrelation goes to exactly zero at the Wendland truncation radius in the time dimension (the posterior median of which is 3.8 days). Importantly, this demonstrates that defining bump functions only in the spatial domain (and not time) in combination with a dense core kernel (the Matérn) results in local non-sparsity. The reference site at Houston, TX is located inside an activated bump function and shares at least one bump function with the four comparison points; therefore, their autocorrelation is again never identically zero. Finally, the reference site near Manhattan, NY is not contained within an activated bump function and hence shares a bump function with none of the four comparison sites. This means that the autocorrelation between all pairs of sites is exactly zero after 3.8 days. Similar behavior emerges for the other four sites of interest.
Figure 7: Directional autocorrelation functions for pairs of locations over time. Here, we focus on three reference locations: Denver, CO; Houston, TX; and Manhattan, NY. For each of these reference locations, we compare correlation at four sites that are $2^{\circ}$ north, south, east, and west. In panel (a), the activated bump functions (i.e., those with $a_{ij}=1$) are shown in dashed gray. In panel (b), correlations are only plotted when they are nonzero.
@@
Next, for the same test sites, we assess the time-evolving behavior of daily maximum temperatures over 2001-2005; see Figure 8 (the first two years are omitted for visibility of results). For this very large data set, taking draws from the posterior predictive distribution described in Section 3.2 is computationally infeasible; instead, we simply calculate the kriging mean and covariance ${\bf m}_{{\bf y}|{\bf z}}$ and ${\bf C}_{{\bf y}|{\bf z}}$ using the posterior median of $\boldsymbol{\beta}$} and $\boldsymbol{\theta}$}. The training sites with test locations are shown on the top left, with the test data and posterior predictions shown in each thumbnail. The most obvious feature of the predictions is the annual cycle, which is well-characterized by our prior mean function. However, note that the specific shape of the annual cycle is flexibly modeled across sites: daily maximum temperatures in Houston, TX (top right) increase sharply in late winter and decrease sharply in late fall, while temperatures in Portland, OR (middle left) increase and decrease much more slowly between the cold and warm season. The GP predictions maintain the large day-to-day variations present in daily maximum temperatures, which is reflected in the data. Also note that the borrowing of strength over space and time allows the GP to learn the time-varying structure even when there are large chunks of missing data in the training data, e.g., in the first half of 2003 at the Houston, TX gauged location. Our spatio-temporal modeling provides a significant benefit in terms of prediction accuracy relative to inverse-distance weighting: as shown in each thumbnail of Figure 8, the improvement in RMSE ranges from 11.5% (at the site near Nashville, TN) to more than 30% (at the site near Manhattan, NY site). Finally, the three-sigma GP uncertainties, represented by the red shaded band in Figure 8, are clearly nonstationary as well: note the larger magnitude of uncertainties at, e.g., Houston, TX compared to the smaller uncertainties for the Berkeley, CA site.